A factory produces
6 motorcycles in 9 hours. Write a
proportion and solve it to find how
many hours it takes to produce
16 motorcycles.

Answer by stanbon(75887)

A factory produces
6 motorcycles in 9 hours. Write a
proportion and solve it to find how
many hours it takes to produce
16 motorcycles.
-----
x/16 = 9/6
x = 16(9/6)
x = 16(3/2)
x = 24 hours to produce 16 motorcycles
===================
Cheers,
Stan H.